ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 246, 2021

Adrian, M. and E. Kaplan On The Langlands parameter of a simple supercuspidal representation: even orthogonal groups

459

Alon, N., S. Ganguly and N. Srivastava High-girth near-Ramanujan graphs with localized eigenvectors

1

Athanasiadis, Ch. A. Face numbers of barycentric subdivisions of cubical complexes

423

Bardadyn, K. and B. K. Kwaśniewski Spectrum of weighted isometries: C^{∗}-algebras, transfer operators and topological pressure

149

Braunling, O., R. Henrard and A.-C. van Roosmalen K-theory of locally compact modules over orders

315

Czudek, K., T. Szarek and H. Wojewódka-Ściążko The law of the iterated logarithm for random interval homeomorphisms

47

d’Elbée, C. and Y. Halevi Dp-Minimal integral domains  487
Frauenfelder, U. Helium and Hamiltonian delay equations  239

de Freitas, T. H., V. H. Jorge Pérez and A. J. Miranda Gluing of analytic space germs, invariants and Watanabe’s conjecture

211

Ganguly, S. See N. Alon Goldstern, M., J. Kellner, D. A. Mejía and S. Shelah

Preservation of splitting families and cardinal characteristics of the continuum

73

Gregory, W. and A. Koldobsky Inequalities for the derivatives of the Radon transform on convex bodies

261

Halevi, Y. See C. d’Elbée

Hamilton, L. and A. Moitra The Paulsen problem made simple  299

Havrilla, A. and T. Tkocz Sharp Khinchin-type inequalities for symmetric discrete uniform random variables

281

Henrard, R. See O. Braunling Jorge Pérez, V. H. See T. H. de Freitas Kaplan, E. See M. Adrian Kellner, J. See M. Goldstern Koldobsky, A. See W. Gregory Kwaśniewski, B. K. See K. Bardadyn Lapid, E. On an inequality of Bushnell–Henniart for Rankin–Selberg

conductors 131 Lev, N. and G. Reti Poisson summation formulas involving the sum-

of-squares function

403

Lipham, D. S. A characterization of Erdős space factors  395

Mejía, D. A. See M. Goldstern Miranda, A. J. See T. H. de Freitas Moitra, A. See L. Hamilton Montgomery, R., A. Pokrovskiy and B. Sudakov C_{4}-free sub-

graphs with large average degree

55

Morotti, L. and H. P. Tong-Viet Proportions of vanishing elements in finite groups

441

Olevskii, A. and A. Ulanovskii On the Szegő–Kolmogorov prediction theorem

335

Pitz, M. Proof of Halin’s normal spanning tree conjecture  353

Pokrovskiy, A. See R. Montgomery Reti, G. See N. Lev van Roosmalen, A.-C. See O. Braunling Sampaio, J. E. Multiplicity, regularity and Lipschitz geometry of real

analytic hypersurfaces

371

Shelah, S. Isomorphic limit ultrapowers for infinitary logic  21

Shelah, S. See M. Goldstern Srivastava, N. See N. Alon Sudakov, B. See R. Montgomery Szarek, T. See K. Czudek Tkocz, T. See A. Havrilla Tong-Viet, H. P. See L. Morotti Ulanovskii, A. See A. Olevskii Wojewódka-Ściążko, H. See K. Czudek